​(Stock repurchase and​ taxes)  The Barryman Drilling Company is planning on repurchasing ​$1.13 million worth of the​ company's 500,000 shares of​ stock, which is currently trading at a price of ​$11.25 per share. Stan Barryman is the founder of the company and still holds 15,000 shares of company stock that he originally purchased for ​$7.97 per share. If Stan decides to sell 2,000 of his shares for ​$11.25 a​ share, what will be his​ after-tax proceeds where capital gains are taxed at 15 ​percent?

​Stan's after-tax proceeds from the sale are $ nothing. ​(Round to the nearest​ dollar.)

Tax on gain:

After tax proceed:
Answer #1

​1: Stan's after-tax proceeds from the sale are

= (Proceeds from sale – Cost)*(1-Tax rate)

= (2000*11.25-2000*7.97)*(1-0.15)

=5576

2: Tax on gain:

(Proceeds from sale – Cost)*Tax rate

= (2000*11.25-2000*7.97)*0.15

=984

3: After tax proceed:

=(Proceeds from sale – Cost)*(1-Tax rate)

= (2000*11.25-2000*7.97)*(1-0.15)

=5576